9. Find the change of coordinates matrix P from the basis B = {1 + 2t, 2 + 3t} to the basis C = {t, 1 + 5t} of P1.
Answer #1

Here the given 91 bases are and B = { it af, 2+3+ } e-{+, 1+ 5t} of Pi Now, t = a (1+2+) at b (2+3t) t = a +zat +26+obt er, tit= 2 (1+26) - (2+3+) Again, 1 +5t (1426) + 2 (2+ 3+) or, 1+5t - cezettad + o2t 1+52 = (+2) + t (2e esd) which gives (3) @f2dItst = 4 (1+2+)-(2+3+) Therefore, t= 2(1+2+) - (2+36) 1+5+ = 4(1424 ) - 3(2+36) The Change of coordinates makoix p to the bas